what are 3 ordered pairs that are solutions to 3x+y=1
Answer by rfer(16322)

3X+Y=1
0+Y=1
3X+0=1
3X=1
X=1/3
3(2)+Y=1
6+Y=1
Y=-5
(0,1)
(1/3,0)
(2,-5)